1. Artis Sales has two store locations. Store A has fixed costs of $210,000 per month and a variable cost ratio of 60%. Store B has fixed costs of $390,000 per month and a variable cost ratio of 30%. What is the break-even sales volume for Store A?

2.  Liu Sales has two store locations. Sanford has fixed costs of $169,000 per month and a contribution margin ratio of 30%. Orlando has fixed costs of $400,000 per month and a contribution margin ratio of 70%. At what sales volume would the two stores have equal profits or losses?

3.

Travon's Limo Service provides transportation services in and around Bentonville. Its profits have been declining, and management is planning to add a package delivery service that is expected to increase revenue by $275,000 per year. The total cost to lease additional delivery vehicles from the local dealer is $60,000 per year. The present manager will continue to supervise all services. However, labor and utilities costs will increase by 40% and rent and other costs will increase by 15% when the package delivery service is added.

Travon’s Limo Service
Annual Income Statement
Before Expansion
Sales Revenue $ 960,000
Costs:
Vehicle leases $ 400,000
Labor 290,000
Utilities 50,000
Rent 100,000
Other Costs 60,000
Manager’s Salary 120,000
Total Costs 1,020,000
Operating Profit (Loss) $ (60,000 )



a. Prepare a report of the differential costs and revenues if the delivery service is added.
b. Should management start up the delivery service? Explain your answer.
Answer #1

Question 1

Break Even Sales Volume for Store A = Fixed Costs / Contribution Margin Ratio

Fixed Costs for Store A = $ 210,000

Contribution Margin Ratio = 1 - Variable Cost Ratio

Variable Cost Ratio = 60%

Contribution Margin Ratio = 1 - 60%

Contribution Margin Ratio = 40%

Break Even Volume for Store A = 210,000 / 40%

Break Even Volume for Store A = $ 525,000

Question 2

Let Sales Revenue for both the Stores = $ X

Sales Revenue of Sanford * Contribution Margin Ratio of Sanford - Fixed Costs of Sanford = Sales Revenue of Orlando * Contribution Margin Ratio of Orlando - Fixed Costs of Orlando

X * 30% - 169,000 = X * 70% - 400,000

0.30X - 169,000 = 0.70X - 400,000

400,000 - 169,000 = 0.70X - 0.30X

231,000 = 0.40X

X = 231,000 / 0.40

X = $ 577,500

At the Sales Revenue of $ 577,500 Both Orlando and Sanford Stores will have a equal Profit of $ 4,250

For Sanford

Profit at Sales of 577,500 = 577,500 * 30% - 169,000 = $ 4,250

For Orlando

Profit at Sales of 577,500 = 577,500 * 70% - 400,000. = $ 4,250

Question 3

Differential Report Analysis

A B C=A-B
Particulars Before Expansion After Expansion Difference
Sales Revenue 960,000 12,35,000 275,000
Costs:
Vehicle Leases 400,000 460,000 60,000
Labor 290,000 406,000 116,000
Utilities 50,000 70,000 20,000
Rent 100,000 115,000 15,000
Other Costs 60,000 69,000 9,000
Manager Salary 120,000 120,000 -
Total Costs 10,20,000 12,40,000 220,000
Operating Profit (Loss) (60,000) (5,000) 55,000

Sales Revenue after Expansion = 960,000 + 275,000 = $ 12,35,000

Vehicle Leases after Expansion = 400,000 + 60,000 = $ 460,000

Labour Costs after Expansion = 290,000 * 140% = $ 406,000

Utilities Costs after Expansion = 50,000 * 140% = $ 70,000

Rent Cost after Expansion = 100,000 * 115% = $ 115,000

Other Costs after Expansion = 60,000 * 115% = $ 69,000

Requirement B

Yes the delivery service should be added as it would reduce the level of operating loss for the company as a whole.